if mC2=n , then find the value of (m+1)C4
Answer by stanbon(75887)

if mC2=n , then find the value of (m+1)C4
----
mC2 = [m(m-1)/2] = n
So, m(m-1) = 2n
-----
(m+1)C4 = (m+1)(m)(m-1)(m-2)/(1*2*3*4)
------
Substitute to get:
(m+1)(2n)(m-2) = (m+1)C4
=============================
Cheers,
Stan H.